# Federated Unlearning: A Survey on Methods, Design Guidelines, and Evaluation Metrics

### v1

@@ III-B Why Is Unlearning Needed if Data Are Not Shared?
To intuitively understand the impact of a client’s data on the global model we performed a set of experiments, whose results are shown in Table III. Specifically, we compared the performance in accuracy of a global model trained including or excluding a specific client $u$ in the federation. While the test accuracy is very similar among such two versions of the global model, the global model exhibits significantly higher accuracy on client $u$’s train data when the federation includes client $u$. This means that the global model trained with client $u$ among the federation may leak information about client $u$’s private data and that may be susceptible to, at least, membership attacks [^2] As we will deepen in Sec. IV-B, membership attacks tries to determine whether one or more specific samples have been used for training.. It is worth noting that the same trend emerges with homogeneous and heterogeneous data distributions among clients.
Table III reports the accuracy of a global model $w$ trained on a federation of participants including or excluding a specific client $u$. Three datasets for image classification were considered (CIFAR-100, Birds and Aircraft), and a visual transformer architecture was used (MiT-B0 [33]), starting from a pretrained checkpoint and performing FL for 100 rounds. The test column refers to the global model’s accuracy on a test set that clients never see during training. The train column refers to the global model’s accuracy on the train set of client $u$. We tracked the accuracy of the global model on test and client $u$’s train data across 100 rounds when the federation includes client $u$, visually highlighting whether client $u$ participates in a round. As evident from Figure 4, when client $u$ is included in the training round, the global model spikes up to a higher level of accuracy on client $u$’s train set. Also, it is worth noting that if client $u$ does not participate for a while, the global model tends to become less accurate on that train data, but still above the test performances. This phenomenon has also been noted in [26].
Moreover, to understand whether natural forgetting could be a viable option for unlearning, i.e. just excluding the client that requested its removal from the federation, we performed a set of further experiments.We want to empirically determine whether the contribution of that client naturally fades away after one or a few rounds, taking advantage of catastrophic forgetting (see also Sec. II-C). We simulated a large-scale federation with 500 clients training for 100 rounds (4% participation rate), with client $u$ leaving at round 50. As depicted in Fig. 5a (homogeneous data among clients), after client $u$ leaves and for the rest of the FL process, the global model steadily keeps generating more meaningful predictions than the retrained baseline on client $u$’s train data for all the tracked rounds. Furthermore, the global model still outputs more accurate predictions (with lower loss as well) on client $u$’s train data compared to a sanitized federation from the beginning (see also the complete metrics in Fig. 8). Contrarily, when data are heterogeneous (reported in Fig. 5b), catastrophic forgetting seems to help natural unlearning: after a few rounds, there is no sensible difference between the tracked metrics for the two versions of the global model.
Overall, the empirical evidence shows that, even with a large-scale federation of clients holding small local datasets, the influence of a client’s contribution on the global model may remain noticeable for long at least when data are homogeneously distributed.
Further details about the experimental settings, employed metrics, and reference to our code repository are provided in Appendix A.
@@
Dataset Test Train Test Train
@@
III-C Challenges in Federated Unlearning
@@
Iterative Learning Process. FL algorithms proceed in iterative rounds, and the global model $w_{t}$ aggregated at round $t$ is the starting state for round $t+1$. Suppose that at round $t_{i}$ a client $u$, which has been included in a round $t_{u}$ ($t_{u}<t_{i}$), requires its contribution deletion (e.g., client unlearning). The naive solution to client unlearning would be to recover the global model at round $t_{u}$ and to exclude client $u$’s updates from the aggregation of round $t_{u}$. However, the iterative nature of FL algorithms implies that all the global aggregations subsequent to round $t_{u}$ should be invalidated.
@@
Non-deterministic Training Process. The selection of clients that participate in an FL round is subject to various dynamics. For example, clients that are eligible for a certain round may not be available in the future, due to connection or battery issues, or due to unwillingness to participate. This peculiarity of FL makes the possibility of exactly reproducing the rounds after $t_{u}$ unrealistic. At the same time, discarding all the knowledge acquired between the aggregation at round $t_{u}$ and the aggregation at round $t_{i}$ (i.e., the most recent round), that possibly may be distant in time with $t_{u}<<t_{i}$, would waste significant learning efforts of the federation. In this sense, methods that can embed that knowledge are most appealing.
@@
Inscrutable Model Updates and Finite Memory. In principle, the FL aggregator or server should not or could not be able to inspect the model updates of single clients (e.g., because the updates could be aggregated via Multi-Party Computation [34, 35]). This would make unfeasible the naive solution of rolling back to the global aggregation of round $t_{u}$ because it requires access to single updates of clients. At the other end of the spectrum, the central FL aggregator should store the entire history of aggregated models and collected updates of each round, which would translate into unfeasible storage capacity.
@@
Inscrutable Local Datasets. While in centralized ML all the data used for training are supposed to be directly accessible and potentially be used for MU. Contrary, in FL contexts, datasets remain private by design. This prevents the application of solutions from plain MU that require direct access to the data to forget and/or to the data to retain. For example, the recent work in [36], targeting centralized unlearning, selectively induces noisy knowledge on the data to forget while reinforcing good knowledge on the retain set respectively employing an incompetent or competent teacher in a Knowledge-Distillation framework. This kind of mechanism cannot be directly adapted to FL due to the inscrutability of local datasets of clients, which cannot be exported or accessed by other parties.
@@
Data Heterogeneity. In centralized ML the data is supposed to be identically and independently distributed (IID) since data are centralized in the same location and can be shuffled and arbitrarily partitioned among machines in case of distributed training. On the contrary, one of the defining characteristics of the FL setting is represented by the presence of data heterogeneity among clients, i.e. clients hold data that are likely to be non-IID (see Fig. 6 in Appendix for an example of non-IID data among clients). While this may hamper the convergence of the global model [37, 38], this trait of FL settings could be convenient for an FU algorithm, since it may be easier to erase the learned representations peculiar to a skewed subset of data (see also the illustrative results in Sec. III-B). However, since data are not directly accessible, it may be not straightforward to assess the level of data heterogeneity in advance.
@@ III-D Entity Performing Unlearning
  • Target Client. The target client, i.e. client $u$, has direct access to the data to be forgotten. It can locally perform and assess the unlearning success. However, this poses further security concerns when considering possibly malicious clients, who may request a local unlearning to inject a backdoor in the model or to hamper the training on purpose, taking advantage of their role in the unlearning phase. Furthermore, regular MU methods often require both retain and forgetting data, e.g., to alternate forgetting and retaining phases so that the model can selectively unlearn. The technical issue in only-local FU is to be able to retain good knowledge without having access to other clients’ data.
  • Remaining Clients: The remaining clients, i.e. the federation without client $u$, hold the retain data and are typically used to recover the performance of an unlearned model.
Given the above benefits and drawbacks, mixed solutions may be also possible, e.g., target client $u$ and the server may cooperate during the unlearning phase.
@@ IV-A Requirements of Federated Unlearning Algorithms
    Improved Efficiency wrt. the Retraining Strategy. An unlearning algorithm should be more efficient than the naive, possibly unfeasible, retraining strategy, e.g. re-utilizing the knowledge that would have been discarded by a global model rollback
@@
    • Comparable performances on test data wrt. the original model. The performance (e.g., loss and accuracy) of the unlearned model on a test set should be comparable to the model’s performance before the unlearning algorithm is applied.
@@ IV-B Metrics to Assess Unlearning
Improved Efficiency Metrics and Indicators. This class of metrics relates to the Requirement (1) of the previous section and usually compares, as a primary dimension, the time needed by the proposed unlearning algorithm to recover performance similar to a baseline, e.g. the original model or the retrained model. For example, the ratio among $T_{w^{*}}/T_{w_{\bar{u}}}$, with $T_{w^{*}}$ being the time needed by a model trained with a specific baseline, e.g. $w^{r}$, and with $T_{w^{\bar{u}}}$ being the time needed by the unlearned model to attain a certain test performance, can be used as an indicator of speed-up. However, recovery time is not the only measure used in the FU literature. In [49], Halimi et al. consider the required cumulative communication cost in client-to-server model updates to achieve similar performance (e.g., in test accuracy) to assess the efficiency of the unlearned method compared to the retrained model, with the lower the better. This comparison could be also expressed as a ratio $Bytes_{w^{*}}/Bytes_{w_{u}}$ (higher is better), with $Bytes$ representing the size (e.g., in MB) of the cumulative client-to-server updates.
In [50], Su et al. also focus on minimizing the number of clients that participate in the recover process.
Retained or Recovered Performance Metrics. This class of metrics relates to the Requirement (2) of the previous section. The straightforward metrics that can be employed to asses whether the unlearned model achieves performances similar to a baseline model are the test performances (e.g., loss, accuracy). Another potential metric to consider is the Expected Calibration Error (ECE) [39]. It is used to evaluate the calibration of probabilistic models, focusing on the alignment between predicted probabilities and actual observed frequencies. It involves dividing predictions into probability bins, comparing average predicted probabilities with observed accuracy in each bin, and calculating the average absolute differences. ECE provides a measure of how well a model’s predicted probabilities reflect the true likelihood of events, indicating the model’s reliability and confidence calibration. In [51], Liu et al. propose to use a compact indicator for comparing the test performances of the unlearned model to a reference model. Such a metric is the Symmetric Absolute Percentage Error (SAPE), computed as
@@
with the baseline model indicated with $w^{*}$ (e.g., the retrained model or the original model) and $Acc({\cdot},D_{test})$ being the accuracy on test data. SAPE spans in a range $\in[0,1]$, and the lower the value, the closer the unlearned model is to match the baseline model’s test performance. To assess Class Unlearning, the test performance should be measured and compared only on the subset of test data belonging to the retained classes [52].
Forgetting Verification Metrics and Indicators. This class of metrics relates to the Requirement (3) of the previous section. Comparing the performance (e.g., loss and accuracy) of the unlearned model to a baseline model (e.g., the retrained model) on client $u$’s train data can represent a preliminary indicator to asses selective forgetting. For example, the retrained model and the unlearned model should exhibit similar performance on client $u$’s train data. However, from a real-world, practical perspective, this can be only evaluated locally by clients, since their data are private.
@@
In [26], Gao et al. propose to employ metrics on specific subsets of client $u$’s forgetting data, called markers, on which calculating more reliable indicators about unlearning performance. Gao et al., most notably, consider as markers the forgettable samples, i.e. samples with the highest loss variance across several rounds, and erroneous samples, i.e. the high loss samples. To assess unlearning, Gao et al. mainly track four metrics: loss, accuracy, influence function (IF) [41], which estimates the impact of a training sample on model prediction, and a metric based on KL divergence. The latter is computed as follows:
@@
The formulation above measures the distributional distance between a model’s output on forgetting data $f_{w}(S_{u})$ and a uniformly distributed output probability $p$, with $p=(\frac{1}{L},...,\frac{1}{L})$, $L$ the number of total classes, and $KL$ the KL divergence. If the KL divergence is low or even 0, the model is not able to produce any meaningful predictions on forgetting data, i.e. the model has forgotten such data. The KL divergence can be also used to measure the distance between weight distributions, i.e. to assess the disparity between the weight distribution obtained through the unlearning process and that obtained through retraining [54], with lower values expected.
While the metrics detailed above are mainly employed to assess the forgetting in client or sample unlearning methods, other metrics tailored to evaluate class unlearning exist. For example, suppose a test set is divided into two subsets, i.e., the set of samples labeled with the unlearned class and the set of all the other samples. The unlearned model’s accuracy on the former set should ideally be 0, whereas the accuracy on the latter set should be similar to the retrained model [55, 52]. However, it should be ensured that the unlearned model does not just misclassify the removed class with another. A possible metric to assess that the unlearned model spans the removed class’s prediction across the remaining classes is presented in [52]. Wang et al. [52] measures the KL divergence among the distribution of the unlearned model’s per-class accuracy and the distribution of retrained model’s per-class accuracy. The closer to 0 the distance, the more similar the distributions are.
@@ V Literature Review of Federated Unlearning Methods
This section offers a comprehensive review of the existing literature on FU methods. We categorize the referenced works based on the unlearning objectives, techniques, and metrics employed for evaluation. The key features of the reviewed papers are summarized in Table V. Each referenced paper is linked with an unlearning objective determined by the respective authors proposing the method. The Where column indicates the entities involved in the unlearning process, while the Metrics column reports the identifiers of the metrics (refer to Table IV) employed for evaluating the effectiveness of the unlearning scheme. Additionally, the Proxy Data column denotes whether supplementary data, typically required by the server for recovery, are necessary. Finally, we highlight if the authors have made their solution accessible to the community or if they plan to do so upon acceptance.
@@
Objective Technique Ref. Where Metrics Proxy Data Open Source Client Re-calibration of Historical Updates [53] Remaining Clients & Server IE.1, RP.2, FV.1, FV.12 - [56] Remaining Clients & Server RP.2, FV.11 - - [57] Remaining Clients & Server IE.2, RP.2, FV.11 - - [26] Target Client & Server IE.1, IE.2, FV.1, FV.3, FV.8 - - [58] Server IE.1, IE.2, RP.2, FV.1 Yes - Knowledge Distillation [59] Server FV.1, FV.11 Yes - [60] Target Client & Neighborhood RP.2 Yes - [55] All Clients & Server IE.1, RP.3, FV.1, FV.11 - - Gradient Modification [49] All Clients & Server IE.1, IE.2, RP.3, FV.11, FV.12, FV.14 - - [61] All Clients & Server RP.2, FV.11 - - [62] Target Client RP.2, FV.11 - Clustering [50] Clients in the Cluster & Server IE.1, RP.2 - Bayesian FL [54] Target Client FV.5 - - [63] Target Client RP.3, FV.1 - - [64] Target Client RP.2, RP.3, RP.5, FV.1 - - [65] Target Client IE.1, RP.2, RP.3, FV.1, FV.5, FV.6, FV.15 - Differential Privacy [66] Server IE.1, RP.2, FV.12 - - Other Approaches [67] Server IE.1, RP.2, RP.3 - - [68] Target Client & Server IE.1, RP.2 - Class Pruning [52] Target Client IE.1, RP.3, FV.1, FV.2, FV.4, FV.12 - - Knowledge Distillation [55] All Clients & Server IE.1, RP.3, FV.1, FV.11 - - Sample Gradient Modification [51] All Clients IE.1, RP.1, RP.2 - [69] Target Clients & Remaining Clients IE.1, IE.2, RP.3, RP.4, FV.1, FV.13 - [70] Target Client & Server IE.1 Yes - [71] Target Client RP.2, RP.3, FV.1, FV.13 Yes - [40] Target Client RP.2, RP.3, FV.7 - - [72] Target Client RP.2, FV.12 - - Quantization [73] Target Client & Server IE1, RP.1, FV.12 - [74] All Clients & Server IE.1, RP.2, RP.3, FV.1 - - Reinforcement Learning [42] Target Client RP.2, FV.9, FV.10 - - Other Approaches [75] All Clients & Server IE.1, IE.2, RP.2, FV.1 - [68] Target Client & Server IE.1, RP.2 -

### v3

@@ III-B Why Is Unlearning Needed if Data Are Not Shared?
To intuitively understand the impact of a client’s data on the global model we performed a set of experiments, whose results are shown in Table III.
Specifically, we compared the performance in accuracy of a global model trained including or excluding a specific client $u$ in the federation. While the test accuracy is very similar among such two versions of the global model, the global model exhibits significantly higher accuracy on client $u$’s train data when the federation includes client $u$. This means that the global model trained with client $u$ among the federation may leak information about client $u$’s private data and that may be susceptible to, at least, membership attacks [^2] As we will deepen in Sec. IV-B, membership attacks tries to determine whether one or more specific samples have been used for training.. The same trend emerges with homogeneous and heterogeneous data distributions among clients.
Table III reports the accuracy of a global model $w$ trained on a federation of participants including or excluding a specific client $u$. Three datasets for image classification were considered (CIFAR-100, Birds and Aircraft), and a visual transformer architecture was used (MiT-B0 [35]), starting from a pretrained checkpoint and performing FL for 100 rounds. The test column refers to the global model’s accuracy on a test set that clients never see during training. The train column refers to the global model’s accuracy on the train set of client $u$. We tracked the accuracy of the global model on test and client $u$’s train data across 100 rounds when the federation includes client $u$, visually highlighting whether client $u$ participates in a round. As evident from Figure 4, when client $u$ is included in the training round, the global model spikes up to a higher level of accuracy on client $u$’s train set. It is worth noting that if client $u$ does not participate for a while, the global model tends to become less accurate on that train data, but still above the test performances. This phenomenon has also been noted in [28].
@@
Dataset Test Client $u$ Train Data Test Client $u$ Train Data
@@
Moreover, to understand whether natural forgetting could be a viable option for unlearning, i.e., just excluding the client that requested its removal from the federation, we performed a set of further experiments.We want to empirically determine whether the contribution of that client naturally fades away after one or a few rounds, taking advantage of catastrophic forgetting (see also Sec. II-C). We simulated a large-scale federation with 500 clients training for 100 rounds (4% participation rate), with client $u$ leaving at round 50. As depicted in Figure 5a (homogeneous data among clients), after client $u$ leaves and for the rest of the FL process, the global model steadily keeps generating more meaningful predictions than the retrained baseline on client $u$’s train data for all the tracked rounds. Furthermore, the global model still outputs more accurate predictions (with lower loss as well) on client $u$’s train data compared to a sanitized federation from the beginning (see also the complete metrics in Figure 6). Contrarily, when data are heterogeneous (reported in Figure 5b), catastrophic forgetting seems to help natural unlearning: after a few rounds, there is no sensible difference between the tracked metrics for the two versions of the global model.
Overall, the empirical evidence shows that, even with a large-scale federation of clients holding small local datasets, the influence of a client’s contribution on the global model may remain noticeable for long at least when data are homogeneously distributed. Further details about the experimental settings, employed metrics, and reference to our code repository are provided in Appendix Appendix - Experimental Setup.
III-C Challenges of Adapting MU in FL Settings
@@
Non-deterministic Training Process. Most of the MU mechanisms suppose that retained data are available when unlearning is performed [6]. Contrarily, the data available in an FL round depends on the client selection, which is subject to various dynamics. For example, clients eligible for a certain round may not be available in the future due to connection or battery issues, or unwillingness to participate. This peculiarity of FL makes the possibility of exactly reproducing the rounds after $t_{u}$ unrealistic. At the same time, discarding all the knowledge acquired between the aggregation at round $t_{u}$ and the aggregation at round $t_{i}$ (i.e., the most recent round), that possibly may be distant in time with $t_{u}<<t_{i}$, would waste significant learning efforts of the federation. In this sense, methods that can embed that knowledge are most appealing.
@@
Inscrutable Model Updates and Finite Memory. In principle, the FL aggregator or server should not or could not be able to inspect the model updates of single clients (e.g., because the updates could be aggregated via Multi-Party Computation [36, 37]). This would make unfeasible the naive solution of rolling back to the global aggregation of round $t_{u}$ and removing the individual contribution of client $u$ because it requires access to single updates of clients. Moreover, even if it could be possible to have direct access to the past model updates, the central FL aggregator would be required to store the entire history of aggregated models and collected updates of each round, which would translate into unfeasible storage capacity.
@@
Inscrutable Local Datasets. In centralized ML all the data used for training are supposed to be directly accessible and potentially be used for MU. On the contrary, in FL contexts, datasets remain private by design. This prevents the application of solutions from plain MU that require direct access to the data to forget and/or to the data to retain. For example, the recent work in [38], targeting centralized unlearning, selectively induces noisy knowledge on the data to forget while reinforcing good knowledge on the retain set respectively employing an incompetent or competent teacher in a Knowledge-Distillation framework. This kind of mechanism cannot be directly adapted to FL due to the inscrutability of clients’ local datasets, which cannot be exported or accessed by other parties.
Iterative Learning Process. An additional factor that differentiates FU from MU is the iterative nature of FL. FL algorithms proceed in cyclical rounds, and the global model $w_{t}$ aggregated at round $t$ is the starting state for round $t+1$. Suppose that at round $t_{i}$ a client $u$, which has been included in a round $t_{u}$ ($t_{u}<t_{i}$), requires its contribution deletion (e.g., client unlearning). The naive solution to client unlearning would be to recover the global model at round $t_{u}$ and to exclude client $u$’s updates from the aggregation of round $t_{u}$. However, the iterative nature of FL algorithms implies that all the global aggregations subsequent to round $t_{u}$ should be invalidated.
@@
Data Heterogeneity. In centralized ML the data is supposed to be identically and independently distributed (IID) since data are centralized in the same location and can be shuffled and arbitrarily partitioned among machines in case of distributed training. On the contrary, one of the defining characteristics of the FL setting is represented by the presence of data heterogeneity among clients, i.e., clients hold data that are likely to be non-IID. While this may hamper the convergence of the global model [39, 40], this trait of FL settings could be convenient for an FU algorithm, since it may be easier to erase the learned representations peculiar to a skewed subset of data (see also the illustrative results in Sec. III-B). However, since data are not directly accessible, it may not be straightforward to assess the level of data heterogeneity in advance.
@@ III-D Entity Performing Unlearning
  • Target Client. The target client, i.e., client $u$, has direct access to the data to be forgotten. It can locally perform and assess the unlearning success. However, this poses further security concerns when considering possibly malicious clients, who may request a local unlearning to inject a backdoor in the model or to hamper the training on purpose, taking advantage of their role in the unlearning phase. Furthermore, regular MU methods often require both retaining and forgetting data, e.g., to alternate forgetting and retaining phases so that the model can selectively unlearn. The technical issue in only-local FU is to be able to retain good knowledge without having access to other clients’ data.
  • Remaining Clients. The remaining clients, i.e., the federation without client $u$, hold the retain data and are typically used to recover the performance of an unlearned model.
Given the above benefits and drawbacks, mixed solutions may also be possible, e.g., target client $u$ and the server may cooperate during the unlearning phase.
@@ IV-A Requirements of Federated Unlearning Algorithms
    Improved Efficiency wrt. the Retraining Strategy. An unlearning algorithm should be more efficient than the naive, possibly unfeasible, retraining strategy, e.g., re-utilizing the knowledge that would have been discarded by a global model rollback
@@
    • Performance results on test data wrt. the original model. The performance of the unlearned model on a test set can be worse or even better than that of the original model depending on the impact of the client $u$’s data. For example, when data among clients are highly heterogeneous and a client holds a prevalent fraction or even all of the samples belonging to a class, the related performance results will likely deteriorate due to the difficulty of generalization for that class. Therefore, comparing the performance of the unlearned model with the original model may not be always informative.
@@ IV-B Metrics to Assess Unlearning
Improved Efficiency Metrics and Indicators. This class of metrics relates to the Requirement (1) of the previous section and usually compares, as a primary dimension, the time needed by the proposed unlearning algorithm to recover performance similar to a baseline, e.g., the original model or the retrained model. For example, the ratio among $T_{w^{*}}/T_{w_{\bar{u}}}$, with $T_{w^{*}}$ being the time needed by a model trained with a specific baseline, e.g., $w^{r}$, and with $T_{w^{\bar{u}}}$ being the time needed by the unlearned model to attain a certain test performance, can be used as an indicator of speed-up. However, recovery time is not the only measure used in the FU literature. In [50], Halimi et al. consider the required cumulative communication cost in client-to-server model updates to achieve similar performance (e.g., in test accuracy) to assess the efficiency of the unlearned method compared to the retrained model, with the lower the better. This comparison could also be expressed as a ratio $Bytes_{w^{*}}/Bytes_{w_{u}}$ (higher is better), with $Bytes$ representing the size (e.g., in MB) of the cumulative client-to-server updates. In [51], Su et al. also focus on minimizing the number of clients that participate in the recovery process.
Retained or Recovered Performance Metrics. This class of metrics relates to the Requirement (2) of the previous section. The straightforward metrics that can be employed to assess whether the unlearned model achieves performances similar to a baseline model are the test performances (e.g., loss, accuracy). Another potential metric to consider is the Expected Calibration Error (ECE) [41]. It is used to evaluate the calibration of probabilistic models, focusing on the alignment between predicted probabilities and actual observed frequencies. It involves dividing predictions into probability bins, comparing average predicted probabilities with observed accuracy in each bin, and calculating the average absolute differences. ECE provides a measure of how well a model’s predicted probabilities reflect the true likelihood of events, indicating the model’s reliability and confidence calibration. In [52], Liu et al. propose to use a compact indicator for comparing the test performances of the unlearned model to a reference model. Such a metric is the Symmetric Absolute Percentage Error (SAPE), computed as
@@
with the baseline model indicated with $w^{*}$ (e.g., the retrained model or the original model) and $Acc({\cdot},D_{test})$ being the accuracy on test data. SAPE spans in a range $\in[0,1]$, and the lower the value, the closer the unlearned model is to matching the baseline model’s test performance. To assess Class Unlearning, the test performance should be measured and compared only on the subset of test data belonging to the retained classes [53].
Forgetting Verification Metrics and Indicators. This class of metrics relates to the Requirement (3) of the previous section. Comparing the performance (e.g., loss and accuracy) of the unlearned model to a baseline model (e.g., the retrained model) on client $u$’s train data can represent a preliminary indicator to assess selective forgetting. For example, the retrained model and the unlearned model should exhibit similar performance on client $u$’s train data. However, from a real-world, practical perspective, this can be only evaluated locally by clients, since their data are private.
@@
In [28], Gao et al. propose to employ metrics on specific subsets of client $u$’s forgetting data, called markers, on which calculating more reliable indicators about unlearning performance. Gao et al., most notably, consider as markers the forgettable samples, i.e., samples with the highest loss variance across several rounds, and erroneous samples, i.e., the high loss samples. To assess unlearning, Gao et al. mainly track four metrics: loss, accuracy, and influence function (IF) [42], which estimates the impact of a training sample on model prediction, and a metric based on KL divergence. The latter is computed as follows:
@@
The formulation above measures the distributional distance between a model’s output on forgetting data $f_{w}(S_{u})$ and a uniformly distributed output probability $p$, with $p=(\frac{1}{L},...,\frac{1}{L})$, $L$ the number of total classes, and $KL$ the KL divergence. If the KL divergence is low or even 0, the model is not able to produce any meaningful predictions on forgetting data, i.e., the model has forgotten such data. The KL divergence can be also used to measure the distance between weight distributions, i.e., to assess the disparity between the weight distribution obtained through the unlearning process and that obtained through retraining [55], with lower values expected.
While the metrics detailed above are mainly employed to assess the forgetting in client or sample unlearning methods, other metrics tailored to evaluate class unlearning exist. For example, suppose a test set is divided into two subsets, i.e., the set of samples labeled with the unlearned class and the set of all the other samples. The unlearned model’s accuracy on the former set should ideally be 0, whereas the accuracy on the latter set should be similar to the retrained model [56, 53]. However, it should be ensured that the unlearned model does not just misclassify the removed class with another. A possible metric to assess that the unlearned model spans the removed class’s prediction across the remaining classes is presented in [53]. Wang et al. [53] measure the KL divergence among the distribution of the unlearned model’s per-class accuracy and the distribution of retrained model’s per-class accuracy. The closer to 0 the distance, the more similar the distributions are.
@@ V Literature Review of Federated Unlearning Methods
This section offers a comprehensive survey of the existing literature on FU methods. We categorize the referenced works based on the unlearning objectives, techniques, and metrics employed for evaluation. The key features of the reviewed papers are summarized in Table V. Each referenced paper is linked with the primary unlearning objective specified by the respective authors. The Where column indicates the entities involved in the unlearning process, while the Metrics column reports the identifiers of the metrics (refer to Table IV) employed for evaluating the effectiveness of the unlearning scheme. Additionally, the Proxy Data column denotes whether supplementary data, typically required by the server for recovery, are necessary. Finally, we highlight whether the authors have made their solution accessible to the community (or if they plan to do so).
@@
Objective Technique Ref. Where Metrics Proxy Data Open Source Client Re-calibration of Historical Updates [54] Remaining Clients & Server IE.1, RP.2, FV.1, FV.12 - [57] Remaining Clients & Server RP.2, FV.11 - - [58] Remaining Clients & Server IE.2, RP.2, FV.11 - - [28] Target Client & Server IE.1, IE.2, FV.1, FV.3, FV.8 - - [59] Server IE.1, IE.2, RP.2, FV.1 Yes - Knowledge Distillation [60] Server FV.1, FV.11 Yes - [61] Target Client & Neighborhood RP.2 Yes - [56] All Clients & Server IE.1, RP.3, FV.1, FV.11 - - Gradient Modification [50] All Clients & Server IE.1, IE.2, RP.3, FV.11, FV.12, FV.14 - - [62] All Clients & Server RP.2, FV.11 - - [63] Target Client RP.2, FV.11 - Clustering [51] Clients in the Cluster & Server IE.1, RP.2 - Bayesian FL [55] Target Client FV.5 - - [64] Target Client RP.3, FV.1 - - [65] Target Client RP.2, RP.3, RP.5, FV.1 - - [66] Target Client IE.1, RP.2, RP.3, FV.1, FV.5, FV.6, FV.15 - Differential Privacy [67] Server IE.1, RP.2, FV.12 - - Other Approaches [68] Server IE.1, RP.2, RP.3 - - [69] Target Client & Server IE.1, RP.2 - [70] All Clients & Server IE.1, RP.2, FV.12 - Class Pruning [53] Target Client IE.1, RP.3, FV.1, FV.2, FV.4, FV.12 - - Knowledge Distillation [56] All Clients & Server IE.1, RP.3, FV.1, FV.11 - - Sample Gradient Modification [52] All Clients IE.1, RP.1, RP.2 - [71] Target Clients & Remaining Clients IE.1, IE.2, RP.3, RP.4, FV.1, FV.13 - [72] Target Client & Server IE.1 Yes - [73] Target Client RP.2, RP.3, FV.1, FV.13 Yes - [10] Target Client RP.2, RP.3, FV.7 - - [74] Target Client RP.2, FV.12 - - Quantization [75] Target Client & Server IE1, RP.1, FV.12 - [76] All Clients & Server IE.1, RP.2, RP.3, FV.1 - - Reinforcement Learning [43] Target Client RP.2, FV.9, FV.10 - - Other Approaches [77] All Clients & Server IE.1, IE.2, RP.2, FV.1 - [69] Target Client & Server IE.1, RP.2 - [70] All Clients & Server IE.1, RP.2, FV.12 -